# Netan Dogra

### v1

@@ 8. $2$-adic Coleman integration
The computation of the local height at $p$ involves calculating Frobenius structures and Hodge filtrations on certain filtered $F$-isocrystals $\mathcal{E}$ on $X$ (see [${}^{+}$19, §4] for a similar discussion in the context of ‘non-generalised’ quadratic Chabauty). Roughly speaking, following Nekovar in [BD21] a local $p$-adic height at $p$ is defined to be a function on certain isomorphism classes of filtered $\phi$-modules (with extra structure). These filtered $\phi$-modules are the fibres of $\mathcal{E}$ at $\mathbb{Z}_{p}$-points of $X$.
@@
    If $f=x^{5}-x$ then a basis of $H^{1}_{\mathrm{dR}}(X/K)$ is given by $\{x^{i}\cdot yx:0\leq i\leq 3\}$.
@@ 9.1. The BMSST curve
discussed in the introduction. The $2$-torsion has Galois group $S_{5}$. The class groups of the extensions $K:=\mathbb{Q}_{f}$ and $L:=\mathbb{Q}_{f}^{(2)}$ (where $f:=2x^{5}-32x+16$) are both trivial. $X$ has good reduction outside the primes 139 and 449, where it has semistable reduction. We deduce that $H^{1}_{f}(G_{\mathbb{Q}},\wedge^{2}J[2])$ may be identified with a subspace of
@@
which has dimension $4$. Recall that this implies $\dim H^{1}_{f}(G_{\mathbb{Q}},\wedge^{2}J[2])\leq 3$. Since the Mordell–Weil rank of $J$ is $3$, to deduce finiteness of $X(\mathbb{Q}_{2})_{2}$ we actually need to determine the crystalline condition (note that even if the Mordell–Weil rank was $2$, we would need this to compute $X(\mathbb{Q}_{2})_{2}$). More precisely, let $\mathfrak{p}$ and $\mathfrak{q}$ denote the unique primes above $2$ in $K$ and $L$ respectively. Then finiteness of $X(\mathbb{Q}_{2})_{2}$ will follow from finding the subspace
@@
where $\mathrm{Im}(U)$ is the image of $U$ in $U^{\prime}:=\mathrm{Ker}(L_{\mathfrak{q}}^{\times}\otimes\mathbb{F}_{2}\stackrel{{\scriptstyle\mathrm{Nm}}}{{\longrightarrow}}K_{\mathfrak{p}}^{\times}\otimes\mathbb{F}_{2})$. To do this, we need to compute the crystalline subspace. We do this using our results on integral crystalline lifts and nonabelian $(x-T)$ maps. More precisely, we first computationally verify that $X$ satisfies the conditions of Lemma 24. Hence, if $b$ and $z$ are $\mathbb{Z}_{2}$-points of $X-\infty$ such that $z-b\in 4\cdot J(\mathbb{Q}_{2})$, then Proposition 2 defines an element of $U^{\prime}$ which is a lift of the class $j_{\Gamma,b}(z)$, and Lemma 24 implies that this lies in $\mathcal{L}$ if the image of $X(\mathbb{Q}_{2})$ in $\mathrm{Sym}^{2}H^{1}(\mathbb{Q}_{2},J[2])$ spans $\mathrm{Sym}^{2}H^{1}_{f}(\mathbb{Q}_{2},J[2])$. This reduces verification of (22) to calculating a basis of $\mathcal{O}_{L}^{\times}$, and calculating whether an element of $L_{\mathfrak{q}}$ is a square. Both of these calculations have been studied in previous work on implementing the Chabauty–Coleman method for hyperelliptic curves (see [Sto01]). We carried out this calculation in magma, searching over $\mathbb{Z}_{2}$ points of $X-\infty$ to determine a basis of $\mathcal{L}$ and verify (22).
Having determined the dimension of $\mathrm{Sel}(U_{2})$, we now to describe how to compute $X(\mathbb{Q}_{2})_{2}$. We have the set $S$ of known rational points
@@
span $\mathrm{Sym}^{2}J(\mathbb{Q})\otimes\mathbb{Q}$ (as usual by $P^{2}$ we mean the class of $(P-\infty)\cdot(P-\infty)$ in $\mathrm{Sym}^{2}J(\mathbb{Q})\otimes\mathbb{Q}$).
For $v=139$ and $449$, a regular semistable model has special fibre an elliptic curve intersecting itself at one point. Hence all $\mathbb{Q}_{v}$-points lie on a common irreducible component, and hence the local generalised height at $v$ of $\mathcal{E}(b,z)$ is zero for all $b,z\in X(\mathbb{Q}_{v})$.
@@
the generalised height of $D$ defines an element of $\mathrm{Ker}(\wedge^{2}H^{1}_{\mathrm{dR}}(X/\mathbb{Q}_{2})\to H^{2}_{\mathrm{dR}}(X/\mathbb{Q}_{2}))^{*}/F^{0}$ which is in the image of $H^{1}_{f}(G_{\mathbb{Q}},\wedge^{2}V/\mathbb{Q}_{p}(1))$. Since $\mathrm{Div}^{0}(S)\otimes\mathbb{Q}$ surjects onto $\mathrm{Sym}^{2}J(\mathbb{Q})\otimes\mathbb{Q}$, the rank of $\mathrm{Ker}(\mathrm{Div}^{0}(S)\to\mathrm{Sym}^{2}J(\mathbb{Q})\otimes\mathbb{Q})$ is two. Using the algorithm for computing the local generalised height at $2$ described above, we can verify that the rank of the image of $\mathrm{Ker}(\mathrm{Div}^{0}(S)\to\mathrm{Sym}^{2}J(\mathbb{Q})\otimes\mathbb{Q})$ in $\mathrm{Ker}(\wedge^{2}H^{1}_{\mathrm{dR}}(X/\mathbb{Q}_{2})\to H^{2}_{\mathrm{dR}}(X/\mathbb{Q}_{2}))^{*}/F^{0}$ under the local generalised height is equal to $2$. Hence we can solve for the equations for $X(\mathbb{Q}_{2})_{2}$ described in Lemma 30. In concrete terms, let $t:X(\mathbb{Q})\to\mathbb{Q}_{2}$ send a point $z$ to the $P_{1}-\infty$ coefficient of $z-\infty$ in $J(\mathbb{Q})$ with respect to the basis $P_{1}-\infty,P_{2}-\infty,P_{3}-\infty$. Let
$\Psi:X(\mathbb{Q}_{p})\to\mathbb{Q}_{2}[T]^{\oplus 3}$
@@
$z\mapsto(\log_{J,1}(z-\infty),\log_{J,2}(z-\infty),T)$
and let $\psi:X(\mathbb{Q})\to\mathbb{Q}_{2}^{3}$ be the map $(\log_{J,1}(z-\infty),\log_{J,2}(z-\infty),t(z))$. Let $\mathrm{Sym}^{2}:\mathbb{Q}_{2}^{\oplus 3}\to\mathbb{Q}_{2}^{\oplus 6}$ be the natural map. Choosing a basis of $\mathrm{Ker}(\wedge^{2}H^{1}_{\mathrm{dR}}(X/\mathbb{Q}_{2})\to H^{2}_{\mathrm{dR}}(X/\mathbb{Q}_{2}))^{*}/F^{0}$, view $\widetilde{h}_{2}$ as a map $X(\mathbb{Q}_{2})\to\mathbb{Q}_{2}^{4}.$ For $z\in X(\mathbb{Q})$, we define $\widetilde{\psi}(z):=\mathrm{Sym}^{2}(\psi(z))\oplus\widetilde{h}_{2}(z),$ and similarly define $\widetilde{\Psi}(z)\in\mathbb{Q}_{2}[T]^{10}$ for $z\in X(\mathbb{Q}_{2})$.
For $z\in X(\mathbb{Q}_{2})$ let $M(z)(T)\in\mathrm{Mat}_{10,9}(\mathbb{Q}_{2}[T])$ denote the matrix $(\widetilde{\psi}(P_{1}),\ldots,\widetilde{\psi}(P_{8}),\widetilde{\Psi}(z))$. Let $F_{1}$ and $F_{2}$ denote the determinants of the submatrices obtained by deleting the $9$th and $10$th rows. If $z\in X(\mathbb{Q})$, then setting $T=t(z)$ we see that the rank of $M(z)(T)$ has rank $8$, and hence $F_{1}$ and $F_{2}$ have a common zero, i.e. the Coleman function obtained by taking the resultant $\mathrm{Res}(F_{1},F_{2})$ vanishes at $z$.
Since the codimension of $\mathrm{Sel}(U_{2})$ in $H^{1}_{f}(G_{\mathbb{Q}_{2}},U_{2})$ is one, in general one would not expect that $X(\mathbb{Q}_{2})$ is equal to $X(\mathbb{Q})$, and indeed we found that $X(\mathbb{Q}_{2})$ contained extra points. Specifically, we found that
@@
(see [${}^{+}$08, §10]). They show in particular that any point of $X(\mathbb{Q})$ must be congruent to the known rational points in the statement of Theorem 1 modulo $2^{8}$ (with respect to a smooth model over $\mathbb{Z}_{2}$). Fortunately, none of the ‘fake rational points’ are congruent to known rational points modulo $2^{7}$.
@@
$\wedge^{2}H^{1}_{\mathrm{dR}}(X/\mathbb{Q}_{2})^{*}/(H^{2}_{\mathrm{dR}}(X/\mathbb{Q}_{2})^{*}+F^{0}).$ This is a consequence of our descent calculation, since their divisors vanish in $\mathrm{Sym}^{2}\mathrm{Jac}(X)(\mathbb{Q}(\sqrt{-3}))^{\operatorname{Gal}(\mathbb{Q}(\sqrt{-3})|\mathbb{Q})}$ and hence their generalised heights are in the image of the two dimensional vector space $H^{1}_{f}(G_{\mathbb{Q}},V_{2}J/\mathbb{Q}_{2}(1))$.

### v2

@@ 8. $2$-adic Coleman integration
The computation of the local height at $p$ involves calculating Frobenius structures and Hodge filtrations on certain filtered $F$-isocrystals $\mathcal{E}$ on $X$ (see [BDM+19, §4] for a similar discussion in the context of ‘non-generalised’ quadratic Chabauty). Roughly speaking, following Nekovar, in [BD21] a local $p$-adic height at $p$ is defined to be a function on certain isomorphism classes of filtered $\phi$-modules (with extra structure). These filtered $\phi$-modules are the fibres of $\mathcal{E}$ at $\mathbb{Z}_{p}$-points of $X$.
@@
    If $f=x^{5}-x$ then a basis of $H^{1}_{\mathrm{dR}}(X/K)$ is given by $\{x^{i}\cdot ydx:0\leq i\leq 3\}$.
@@ 9. Examples
Code for the calculations described in this section can be found at https://github.com/netandog/BKdescent.
@@ 9.1. The BMSST curve
discussed in the introduction. The $2$-torsion has Galois group $S_{5}$. The class groups of the extensions $K:=\mathbb{Q}_{f}$ and $L:=\mathbb{Q}_{f}^{(2)}$ (where $f:=2x^{5}-32x+16$) are both trivial. $X$ has good reduction outside the primes 139 and 449, where it has semistable reduction. We deduce that $H^{1}_{f}(\mathbb{Q},\wedge^{2}J[2])$ may be identified with a subspace of
@@
which has dimension $5$. Recall that this implies $\dim H^{1}_{f}(\mathbb{Q},\wedge^{2}J[2])\leq 4$. Since the Mordell–Weil rank of $J$ is $3$, to deduce finiteness of $X(\mathbb{Q}_{2})_{2}$ we need to verify that $\dim H^{1}_{f}(\mathbb{Q},\wedge^{2}V)\leq 2$ (note that even if the Mordell–Weil rank was $2$, we would need this to compute $X(\mathbb{Q}_{2})_{2}$). For any curve $X$ of genus at least one,
$\dim H^{1}(\mathbb{Q},\wedge^{2}T_{2}J)[2]=\dim H^{0}(\mathbb{Q},\wedge^{2}J[2])\geq 1$
via the inclusion of Galois modules $\mathbb{F}_{2}\hookrightarrow\wedge^{2}J[2]$ dual to the Weil pairing. Hence to prove finiteness of $X(\mathbb{Q}_{2})_{2}$ we must prove $\dim H^{1}_{f}(\mathbb{Q},\wedge^{2}J[2])=3$.
More precisely, let $\mathfrak{p}$ and $\mathfrak{q}$ denote the unique primes above $2$ in $K$ and $L$ respectively. Then finiteness of $X(\mathbb{Q}_{2})_{2}$ will follow from finding the subspace
@@
where $\mathrm{Im}(U)$ is the image of $U$ in $U^{\prime}:=\mathrm{Ker}(L_{\mathfrak{q}}^{\times}\otimes\mathbb{F}_{2}\stackrel{{\scriptstyle\mathrm{Nm}}}{{\longrightarrow}}K_{\mathfrak{p}}^{\times}\otimes\mathbb{F}_{2})$. To do this, we need to compute the crystalline subspace. We do this using our results on integral crystalline lifts and nonabelian $(x-T)$ maps. More precisely, we first computationally verify that $X$ satisfies the conditions of Lemma 24. Hence, if $b$ and $z$ are $\mathbb{Z}_{2}$-points of $X-\{\infty\}$ such that $z-b\in 4\cdot J(\mathbb{Q}_{2})$, then Proposition 2 defines an element of $U^{\prime}$ which is a lift of the class $j_{\Gamma,b}(z)$, and Lemma 24 implies that this lies in $\mathcal{L}$ if the image of $X(\mathbb{Q}_{2})$ in $\mathrm{Sym}^{2}H^{1}(\mathbb{Q}_{2},J[2])$ spans $\mathrm{Sym}^{2}H^{1}_{f}(\mathbb{Q}_{2},J[2])$. This reduces verification of (22) to calculating a basis of $\mathcal{O}_{L}^{\times}$, and calculating whether an element of $L_{\mathfrak{q}}$ is a square. Both of these calculations have been studied in previous work on implementing the Chabauty–Coleman method for hyperelliptic curves (see [Sto01]). We carried out this calculation in magma, searching over $\mathbb{Z}_{2}$ points of $X-\{\infty\}$ to determine a basis of $\mathcal{L}$ and verify (22).
Having determined the dimension of $\mathrm{Sel}(U_{2})$, we now describe how to compute $X(\mathbb{Q}_{2})_{2}$. We have the set $S$ of known rational points
@@
span $\mathrm{Sym}^{2}J(\mathbb{Q})\otimes\mathbb{Q}$ (as usual by $P^{2}$ we mean the class of $(P-\{\infty\})\cdot(P-\{\infty\})$ in $\mathrm{Sym}^{2}J(\mathbb{Q})\otimes\mathbb{Q}$).
For $v=139$ and $449$, a regular semistable model has special fibre an elliptic curve intersecting itself at one point. Hence all $\mathbb{Q}_{v}$-points lie on a common irreducible component, and hence the local generalised height at $v$ of $E(b,z)$ is zero for all $b,z\in X(\mathbb{Q}_{v})$.
@@
the generalised height of $D$ defines an element of $\mathrm{Ker}(\wedge^{2}H^{1}_{\mathrm{dR}}(X/\mathbb{Q}_{2})\to H^{2}_{\mathrm{dR}}(X/\mathbb{Q}_{2}))^{*}/F^{0}$ which is in the image of $H^{1}_{f}(\mathbb{Q},\wedge^{2}V/\mathbb{Q}_{p}(1))$. Since $\mathrm{Div}^{0}(S)\otimes\mathbb{Q}$ surjects onto $\mathrm{Sym}^{2}J(\mathbb{Q})\otimes\mathbb{Q}$, the rank of $\mathrm{Ker}(\mathrm{Div}^{0}(S)\to\mathrm{Sym}^{2}J(\mathbb{Q})\otimes\mathbb{Q})$ is two. Using the algorithm for computing the local generalised height at $2$ described above, we can verify that the rank of the image of $\mathrm{Ker}(\mathrm{Div}^{0}(S)\to\mathrm{Sym}^{2}J(\mathbb{Q})\otimes\mathbb{Q})$ in $\mathrm{Ker}(\wedge^{2}H^{1}_{\mathrm{dR}}(X/\mathbb{Q}_{2})\to H^{2}_{\mathrm{dR}}(X/\mathbb{Q}_{2}))^{*}/F^{0}$ under the local generalised height is equal to $2$. Hence we can solve for the equations for $X(\mathbb{Q}_{2})_{2}$ described in Lemma 30. In concrete terms, let $t:X(\mathbb{Q})\to\mathbb{Q}_{2}$ send a point $z$ to the $P_{1}-\{\infty\}$ coefficient of $z-\{\infty\}$ in $J(\mathbb{Q})$ with respect to the basis $P_{1}-\{\infty\},P_{2}-\{\infty\},P_{3}-\{\infty\}$. For $i=1,2$, let
$\log_{J,i}:J(\mathbb{Q}_{2})\to\mathbb{Q}_{2}$
be the map sending $P$ to $\int_{P}\omega_{i}$ for any $\omega_{1},\omega_{2}$ of $H^{0}(X_{\mathbb{Q}_{2}},\Omega_{X|\mathbb{Q}_{2}})$. Let
$\Psi:X(\mathbb{Q}_{2})\to\mathbb{Q}_{2}[T]^{\oplus 3}$
@@
$z\mapsto(\log_{J,1}(z-\{\infty\}),\log_{J,2}(z-\{\infty\}),T)$
and let $\psi:X(\mathbb{Q})\to\mathbb{Q}_{2}^{3}$ be the map $(\log_{J,1}(z-\{\infty\}),\log_{J,2}(z-\{\infty\}),t(z))$. Let $\mathrm{Sym}^{2}:\mathbb{Q}_{2}^{\oplus 3}\to\mathbb{Q}_{2}^{\oplus 6}$ be the natural map sending an element of $\mathbb{Q}_{p}^{\oplus 3}$ to its symmetric square (with respective to any basis of $\mathrm{Sym}^{2}(\mathbb{Q}_{p}^{\oplus 3})$. Choosing a basis of $\mathrm{Ker}(\wedge^{2}H^{1}_{\mathrm{dR}}(X/\mathbb{Q}_{2})\to H^{2}_{\mathrm{dR}}(X/\mathbb{Q}_{2}))^{*}/F^{0}$, view $\widetilde{h}_{2}$ as a map $X(\mathbb{Q}_{2})\to\mathbb{Q}_{2}^{4}.$ For $z\in X(\mathbb{Q})$, we define $\widetilde{\psi}(z):=\mathrm{Sym}^{2}(\psi(z))\oplus\widetilde{h}_{2}(z),$ and similarly define $\widetilde{\Psi}(z)\in\mathbb{Q}_{2}[T]^{10}$ for $z\in X(\mathbb{Q}_{2})$.
For $z\in X(\mathbb{Q}_{2})$ let $M(z)(T)\in\mathrm{Mat}_{10,9}(\mathbb{Q}_{2}[T])$ denote the matrix $(\widetilde{\psi}(P_{1}),\ldots,\widetilde{\psi}(P_{8}),\widetilde{\Psi}(z))$. Let $F_{1}$ and $F_{2}$ denote the determinants of the submatrices obtained by deleting the $9$th and $10$th rows. If $z\in X(\mathbb{Q})$, then setting $T=t(z)$ we see that the rank of $M(z)(T)$ is $8$, and hence $F_{1}$ and $F_{2}$ have a common zero, i.e. the Coleman function obtained by taking the resultant $\mathrm{Res}(F_{1},F_{2})$ vanishes at $z$.
Since the codimension of $\mathrm{Sel}(U_{2})$ in $H^{1}_{f}(\mathbb{Q}_{2},U_{2})$ is one, in general one would not expect that $X(\mathbb{Q}_{2})_{2}$ is equal to $X(\mathbb{Q})$, and indeed we found that $X(\mathbb{Q}_{2})$ contained extra points. Specifically, we found that
@@
(see [BMS+08, §10]). They show in particular that any point of $X(\mathbb{Q})$ must be congruent to the known rational points in the statement of Theorem 1 modulo $2^{8}$ (with respect to a smooth model over $\mathbb{Z}_{2}$). Fortunately, none of the ‘fake rational points’ are congruent to known rational points modulo $2^{7}$.
@@
$\wedge^{2}H^{1}_{\mathrm{dR}}(X/\mathbb{Q}_{2})^{*}/(H^{2}_{\mathrm{dR}}(X/\mathbb{Q}_{2})^{*}+F^{0})$ (modulo a large power of $2$). This is a consequence of our descent calculation, since their divisors vanish in $\mathrm{Sym}^{2}\mathrm{Jac}(X)(\mathbb{Q}(\sqrt{-3}))^{\operatorname{Gal}(\mathbb{Q}(\sqrt{-3})|\mathbb{Q})}$ and hence their generalised heights are in the image of the two dimensional vector space $H^{1}_{f}(\mathbb{Q},V_{2}J/\mathbb{Q}_{2}(1))$.